I am totally confused, please help.
Problem
By using a ruler, you determine that the distance between two cities on a map is 3.4 inches. According to the map’s scale, 2 inches represents 75 miles. What is the actual distance between the two cities?
Questions.
A. Briefly explain the general procedure used to identify and set up a proportion problem.

B. Write a proportion that correctly represents the problem in the Given.
1. Solve the proportion, showing all relevant work.

C. Explain the reasoning behind the equation you set up (e.g., how did you determine the two ratios in the problem, and how do you know if they are equivalent?).

D. Explain each step used to solve the equation. Your explanation should include the original equation, the equation at each step, and the final equation with the unknown equal to some number. (Use complete sentences, not mathematical symbols, to explain the steps.)

E. Briefly explain (in 1–2 sentences) what the answer means in terms of the original problem.





Answer by stanbon(48558)

By using a ruler, you determine that the distance between two cities on a map is 3.4 inches. According to the map’s scale, 2 inches represents 75 miles. What is the actual distance between the two cities?
Questions.
A. Briefly explain the general procedure used to identify and set up a proportion problem.
B. Write a proportion that correctly represents the problem in the Given.
1. Solve the proportion, showing all relevant work.
75/2 = x/3.4
x = 3.4(75/2) = 127.5 miles
C. Explain the reasoning behind the equation you set up (e.g., how did you determine the two ratios in the problem, and how do you know if they are equivalent?).
Your were told 75 actual miles = 2 map inches
You want to know "#of actual miles" = 3.4 inches
D. Explain each step used to solve the equation. Your explanation should include the original equation, the equation at each step, and the final equation with the unknown equal to some number. (Use complete sentences, not mathematical symbols, to explain the steps.)
E. Briefly explain (in 1–2 sentences) what the answer means in terms of the original problem
3.4 inches on the map means 127.5 miles in actuality.
==========
Cheers,
Stan H.